Find the largest of the four even whole numbers whose sum is 100.
Answer by checkley77(7072)

Let x, x+2, x+4, x+6 be the 4 numbers.
x+x+2+x+4+x+6=100
4x=100-2-4-6
4x=88
x=88/4
x=22 ans. for the first number.
22=2=24 ans. for the second number.
22+4=26 ans. for the third number.
22+6=28 ans. for the last number.
22+24+26+28=100
100=100